OpenStudy (anonymous):

Rewrite -54 degrees in radians as a mulitple of pi?

OpenStudy (anonymous):

rad = (degree) pi/180 use this formula.

OpenStudy (mathstudent55):

There are 2 pi radians in a full circle, or 360 deg. That means there are pi radians in 180 degrees. \(\pi ~rad = 180^o\) \(1~rad = \dfrac{180^o}{\pi} \) Find what fraction of 180 deg -54 deg is, and tehn just replace 180 deg with pi radians.